# Multicopy Reinforcement Learning Agents

### v2

@@ 9 Experiments
Figure 3: Varying Costs and Noise. Return and policy during testing for various cost and noise settings, showing the improvement in return from the Joint Action algorithm to Multicopy (a). Actions shown in (b) are best actions. 50 trial runs.
9.1 Learning Curves
In Figure 2(b-c) and (e-f), we compare Multicopy and Joint Action algorithms with various maximum action settings. Figures (b) and (e) use the 7500 episode training time that we will use throughout the remainder of the paper, while Figures (c) and (f) use a longer training time of 50,000 episodes. The Multicopy algorithm does a fairly good job of learning as well in 7500 episodes as it does in 50,000, even with fairly high noise levels. The Joint Action algorithm, however, due to its fractured state space, does not learn as well when training time is short, particularly in (e), where noise is high.
@@ 9.1 Learning Curves
Figure 4: Identical Duplicate Actions. Return and policy during testing for various cost and noise settings when identical duplicate actions are allowed. 50 trial runs.
9.2 Varying noise and costs
As shown in Figure 3(a), the Multicopy algorithm generally performs better than the Joint Actions algorithm in these experiments. The highest proportional improvement is seen when noise is high (0.4) and costs are moderate (-4), and the lowest improvement when noise is 0.0.
Figure 3(b) details the policy chosen in the start state by each algorithm. The Joint Action algorithm is much more likely to choose a simple policy with 1 or 2 actions, perhaps because at this stage it has a better policy for getting across the bridges under those conditions.
The Multicopy algorithm has a more complex pattern of action choices. Generally one would expect that as cost magnitude increases, the agent should make do with fewer copies. Noise should have roughly the opposite effect: as noise increases, the agent should send more copies.
This rule of thumb largely applies for the upper and right rows of the graphs in Figure 3(b). When there is no noise, one action is consistently sufficient (first row). The area where three actions are best forms a wedge from the right side of the table, where costs are low, with a point at about -4, 0.2, where (A, B) and (A, B, C) have roughly the same value.
The number of actions goes back down to one when noise and costs are both high (the extreme lower left region). In this case the agent cannot reliably achieve the task even using all three bridges, and largely returns to using a single copy to keep costs down.
Between the regions where 3 actions are best, and the regions where 1 action is best, there are bands in which 2 actions is the optimal policy.
9.3 Duplicate Action Choices
@@
Identical Actions
When we remove the requirement that there be only one agent per bridge in Figure 4, we see that the agent tends to choose to send multiple agents over the best bridge. In this case is also fairly noisy in its outcome due to its narrowness, so it also fulfills the criteria of Section 5. There is no correlation between the success or failure of the first agent and the second, so making multiple attempts at the same bridge can improve performance even if the agents do no coordination or cooperation on the bridge.
Multiple copies still does not help when there is no noise in the environment (the first rows of all heatmaps in Figure 4). While the Joint Action algorithm is closer in return to the Multicopy algorithm in these experiments (Figure 4(a)), it does make the odd choice to send three agent copies in this case. Generally the "best" action is not consistent across trial runs for this algorithm, however (see Figure 4(b)), indicating that it has not fully converged by this point.
For the Multicopy algorithm, Figure 4(b) shows fewer cases in which the agent chooses fewer than three copies of itself, though there are still a few cases where the agent chooses fewer copies to reduce costs.
9.4 Correlated Dangers
Thus far, we have examined a case in which the agent may send only one copy per bridge and a case in which the agent can and does send multiple copies on the same bridge. Next we examine a case in which despite having the option to send multiple copies on the same path, the rational behavior is to spread copies over different paths. One such scenario occurs when the noise or dangers on a path are correlated for all agents on the same path.
@@ 9.4 Correlated Dangers
Figure 5: Broken Bridges Experiments. Return and policy during testing for various cost and noise settings when identical duplicate actions are allowed and a random bridge is broken in each trial. 20 trial runs.
@@
Since sending multiple agents on the same bridge can lead to both agents being lost, and the Multicopy agent, on the other hand, prefers to spread the copies over at least two different bridges to ameliorate risk, as shown in Figure 5(b).
10 Future Work
There are several research questions that arise from this work.
Wireless Networks
We will continue developing these algorithms in combination with modern function approximation methods such as Neural Networks [1, 15] in order to apply them to Mobile Wireless Networks [13, 10].
Other types of Optimization
In this paper, and in the Mobile Wireless Networks applications that inspired it, the agent seeks to optimize the amount of time required to complete a task. However, other types of optimization functions are possible within the same mathematical framework: building the best object in an assembly line, or getting the farthest in an exploration or rescue task.
Agent communication and coordination
In this work, we assume that communication between agents is expensive or unavailable much of the time, aside from the point at which the agent decides to duplicate itself. We therefore use a model where communication is limited to only the duplication step. This is essential in producing the simplified value function we use here. However, if possible, we will look for types of communication that are possible while maintaining this value function model.
Multiple goals
In some cases, an set of agents might need to achieve several similar goals at once (putting out several fires or finding several items). Part of our ongoing research includes adapting our algorithm to this case.
11 Conclusion
This paper makes several contributions. First, we have shown that for the specific multiagent problem with examine here, considering the joint state space of all agent copies is not necessary, and joint actions are necessary only when duplicating agents.
Second, we have shown that even using a simple very gridworld, the algorithm found a complex pattern of optimization that is easier to construct algorithmically than by hand-coding heuristics.

### v3

@@ 9.1. Learning Curves
Figure 4. Varying Noise and Cost. Return and policy during testing for various cost and noise settings, showing the improvement in return from the Joint Action algorithm to Multicopy (a). Actions shown in (b) are best actions. 50 trial runs.
9.2. Varying noise and costs
As shown in Figure 4(a), the Multicopy algorithm generally performs better than the Joint Actions algorithm in these experiments. The highest proportional improvement is seen when noise is high (0.4) and costs are moderate (-4), and the lowest improvement when noise is 0.0. Figure 4(b) details the policy chosen in the start state by each algorithm. The Joint Action algorithm is much more likely to choose a simple policy with 1 or 2 actions, perhaps because at this stage it has a better policy for getting across the bridges under those conditions.
@@ 9.2. Varying noise and costs
Figure 5. Identical Duplicate Actions. Return and policy during testing for various cost and noise settings when identical duplicate actions are allowed. 50 trial runs.
The Multicopy algorithm has a more complex pattern of action choices. Generally one would expect that as cost magnitude increases, the agent should make do with fewer copies. Noise should have roughly the opposite effect: as noise increases, the agent should send more copies. This rule of thumb largely applies for the upper and right rows of the graphs in Figure 4(b). When there is no noise, one action is consistently sufficient (first row). The area where three actions are best forms a wedge from the right side of the table, where costs are low, with a point at about -4, 0.2, where (A, B) and (A, B, C) have roughly the same value.
The number of actions goes back down to one when noise and costs are both high (the extreme lower left region). In this case the agent cannot reliably achieve the task even using all three bridges, and largely returns to using a single copy to keep costs down. Between the regions where 3 actions are best, and the regions where 1 action is best, there are bands in which 2 actions is the optimal policy.
@@
9.3. Duplicate Action Choices
Identical Actions. When we remove the requirement that there be only one agent per bridge in Figure 5, we see that the agent tends to choose to send multiple agents over bridge A, the shortest bridge (Figure 5(b) Multicopy). This bridge is dangerous due to its narrowness, and there is no correlation between the success or failure between agent copies crossing the bridge. This means that making multiple identical attempts at the same bridge can improve performance even if the agents do no coordination or cooperation on the bridge.
For our Multicopy algorithm, Figure 5(b) shows largely cases in which the agent chooses the maximum 3 copies, though there are still a few cases where the agent chooses fewer copies to reduce costs. Multiple copies also do not help when there is no noise in the environment (the first row of all heatmaps in Figure 5). While the Joint Action algorithm is closer in return to the Multicopy algorithm in these experiments (Figure 5(a)), it does make the odd choice to send three agent copies even when there is no noise. Generally the ”best” action is not consistent across trial runs for this algorithm, however (see Figure 5(b)), indicating that it has not fully converged.
@@ 9.3. Duplicate Action Choices
Figure 6. Broken Bridges Experiments. Return and policy during testing for various cost and noise settings when identical duplicate actions are allowed and a random bridge is broken in each trial. 20 trial runs.
9.4. Correlated Dangers
Thus far, we have examined a case in which the agent is allowed to send only one copy per bridge and a case in which the agent can and does send multiple copies on the same bridge. Next we examine a case in which despite having the option to send multiple copies on the same path, the rational behavior is to spread copies over different paths. One such scenario occurs when the noise or dangers on a path are correlated for all agents on the same path.
@@ 9.4. Correlated Dangers
Since sending multiple agents on the same bridge can lead to both agents being lost, the Multicopy agent prefers to spread the copies over at least two different bridges to ameliorate risk, as shown in Figure 6(b).
10. Future Work
We will continue developing these algorithms in combination with modern function approximation methods such as Neural Networks Arulkumaran et al. (2017); Wong et al. (2023) in order to apply them to Mobile Wireless Networks Tie et al. (2011); Spyropoulos et al. (2008); Manfredi et al. ([n.d.]).
In this paper, and in the Mobile Wireless Networks applications that inspired it, the agent seeks to optimize the amount of time required to complete a task. However, other types of optimization functions are possible within the same mathematical framework: building the best object, or getting the farthest in an exploration or rescue task.
In this work, we assume that communication between agents is expensive or unavailable much of the time, aside from the point at which the agent decides to duplicate itself. We therefore use a model where communication is limited to only the duplication step. This was instrumental in producing the simplified value function we use here. However, we will look for types of communication that are possible while maintaining this value function model. Once there is some level of interaction between agents, non-deterministic policies may be required Vieillard et al. ([n.d.]); Haarnoja et al. ([n.d.]).
Once we move to larger problems and approximate solutions, distributional RL Bellemare et al. ([n.d.]) may prove useful, given a maximization operator for their particle based value representation.
In this paper we primarily looked at different cost/optimization reward ratios in order to show that different multiactions are chosen at different settings. However, even in the single agent case, multiobjective RL Mossalam et al. ([n.d.]); Moffaert and Nowé ([n.d.]); Nguyen et al. ([n.d.]); Hayes ([n.d.]) would be a useful tool to find a range of policies for different weights. Lexicographic ordered RL Tercan and Prabhu ([n.d.]); Hayes et al. ([n.d.])could be particularly useful if it allows us to find the best delay given a particular target for throughput, for example.
11. Conclusion
This paper makes several contributions. First, we have shown that for the specific multiagent problem we examine here, considering the joint state space of all agent copies is not necessary, and joint actions are necessary only when duplicating agents.
Even using a simple very gridworld, the algorithm found a complex pattern of optimization that is easier to construct algorithmically than by hand-coding heuristics.
@@ 11. Conclusion
{acks}
This material is based upon work supported by the National Science Foundation under Grant No. 2154190. Any opinions, findings, and conclusions or recommendations expressed in this material are those of the author(s) and do not necessarily reflect the views of the National Science Foundation.